1. a) Substitute u = sin(x) to evaluate
sin^2(x) cos^3(x) dx.
[trig identity sin2(x)+cos2(x) = 1].
b) Find the antiderivatives:
i) sin(2x) dx
ii) (cos(4x)+3x^2) dx
Answer #1

0 0 (a) () let given that I= l sin? (2) Cos? (N) da sinlay=u cosin Loftsson L1 Substitute - (of(a) dx =du = 1-42 & we have co(in let I I (cos(4x) + 37² ) da = Cop (43) On + 34? da let u=42 de = 4 da o de da = s cofLUS dient + 3.6x2 dr = $s We L) du +